1. Break-even analysis

To be profitable, a firm must recover its costs. These costs include both its fixed and its variable costs. One way that a firm evaluates at what stage it would recover the invested costs is to calculate how many units or how much in dollar sales is necessary for the firm to earn a profit.

Consider the case of Blue Mouse Manufacturers:

Blue Mouse Manufacturers is considering a project that will have fixed costs of $10,000,000. The product will be sold for $41.50 per unit, and will incur a variable cost of $10.75 per unit.

Given Blue Mouse’s cost structure, it will have to sell   units to break even on this project (QBEQBE).

Blue Mouse’s marketing and sales director doesn’t think that the firm’s market is big enough for the firm to break even. In fact, she believes that the firm will be able to sell only about 150,000 units. However, she also thinks that the demand for Blue Mouse’s product is relatively inelastic (so the firm can increase the sales price without significantly decreasing the volume of product sold). Assuming that the firm can sell 150,000 units, what price must it set to break even?

$77.42 per unit

$92.90 per unit

$73.55 per unit

$85.16 per unit

What affects the firm’s operating break-even point?

Several factors affect a firm’s operating break-even point. Based on the scenarios described in the following table, indicate whether these factors would increase, decrease, or leave unchanged a firm’s break-even quantity—assuming that only the listed factor changes and all other relevant factors remain constant.

Increase

Decrease

No Change

The variable cost per unit decreases.
The firm’s tax rate increases.
The firm’s fixed costs increase.

When a large percentage of a firm’s costs are fixed, the firm is said to have a   degree of operating leverage.
Answer #1

If the firm can sell 150000 units at a price of $ x per unit

The total Variable Costs = 10.75*150,000 = $1,612,500

Total Sales = $ x *150000 = 150000 x

At Break Even point , Profit = Sales - Fixed Cost - Variable Cost = 0

So, 150000x- 1612500 - 10000000 = 0

x = 16,125,000/150000 = $77.42 per unit

Hence, to break even the price must be $77.42 per unit

The firm's break-even point is affected by Contribution margin =Sales price per unit - Variable cost per unit as well as Fixed Costs

If the Variable cost per unit decreases, the total variable cost decreases and the firm can break even at a lesser qauantity

Hence , the firm's break even quantity also decreases with decrease in variable cost per unit

If the firm's tax rate increases , there is no change in the break even quantity (as there are no taxes at the operating level)

If the firm's fixed costs increases , the firm has to sell more quantity to break even.

Hence , the firm's break even quantity increases with increase in fixed cost.

When a large percentage of a firm's costs are fixed, the firm is said to have a High  degree of Operating leverage.